# Unraveling Zeroth-Order Optimization through the Lens of Low-Dimensional Structured Perturbations

### v1

@@ 3.3 Towards Improved Convergence: Loss Perspective
    $\mathrm{intdim}\big{(}\mathbf{H}(\bm{\theta}_{t})\big{)}\leq r$ for some constant $r=\Theta(1)$ with respect to $d$.
Remark.  Note that our Assumption 3.8 not only allow for strictly positive $\mu>0$, but also require much smaller local region (i.e., $\|\bm{\theta}-\bm{\theta}_{t}\|$) around the current iterate $\bm{\theta}_{t}$. In Malladi et al. (2023), the authors consider $\mathcal{O}(d)$-ball around $\bm{\theta}_{t}$, which is quite stringent (actually almost global condition), for extremely large $d$ (ex. fine-tuning LLMs). In contrast, our Assumption 3.8 only requires $\mathcal{O}(s)$-ball around $\bm{\theta}_{t}$ since $\mu\sqrt{d}$ is negligible under $\mu=\mathcal{O}(1/\sqrt{dT})$ (refer to Theorem 3.6 or Theorem 3.10).Thus, our Assumption 3.8 significantly alleviate (Malladi et al., 2023) and can really only consider the local behavior of Hessian.
Since we consider both the transformation $\mathbf{M}_{t}$ and the local Hessian $\mathbf{H}(\bm{\theta}_{t})$, we introduce an important quantity for our analysis, the effective overlap, which measures the level of overlapping between two spaces defined by $\mathbf{M}_{t}$ and $\mathbf{H}(\bm{\theta}_{t})$.
Definition 3.9 (Effective Overlap).
For the matrix $\mathbf{M}_{t}$ in (2) and $\mathbf{H}(\bm{\theta}_{t})$ in Assumption 3.8, we define the effective overlap at time $t$ as
$\rho_{t}=\frac{\mathrm{Tr}(\mathbf{M}_{t}\mathbf{H}(\bm{\theta}_{t})\mathbf{M}_{t})}{\sigma_{\text{max}}(\mathbf{M}_{t})^{2}\lambda_{\text{max}}(\mathbf{H}(\bm{\theta}_{t}))}.$ (4)
@@
Intuitively, $\mathrm{Tr}(\mathbf{M}_{t}\mathbf{H}(\bm{\theta}_{t})\mathbf{M}_{t})$ captures “how much the primary directions” of $\mathbf{M}_{t}$ and $\mathbf{H}(\bm{\theta}_{t})$ align and reinforce each other, since both matrices are assumed to be PSD and thus act as “weights” on overlapping subspaces. The normalization by $\sigma_{\text{max}}(\mathbf{M})^{2}\lambda_{\text{max}}(\mathbf{H}(\bm{\theta}_{t}))$ prevents $\rho_{t}$ from inflating merely due to scale, ensuring $\rho_{t}$ remains scale-invariant. If $\mathrm{srank}(\mathbf{M}_{t})=s\ll d$ and $\mathrm{intdim}\big{(}\mathbf{H}(\bm{\theta}_{t})\big{)}=r\ll d$ hold, a non-trivial value of $\rho_{t}$ indicates a significant overlap in those low-dimensional subspaces, which is crucial for effective optimization in high-dimensional spaces. Note also that $0\leq\rho_{t}\leq\min\{r,s\}$ always holds.
We now guarantee the dimension-free convergence.
Theorem 3.10 (Dimension-Free Convergence of ZO-SGD).
Suppose that the same conditions in Theorem 3.6 and Assumption 3.8 hold. Under the following parameter settings
[…]
@@
We make remarks on the convergence results.
@@
3.4 Low-Dimensional Constraints Also Matter for Generalization
All the aforementioned analysis are concerned about the convergence of the zeroth-order SGD. In this section, we aim to provide the generalization error bound of zeroth-order SGD. In pursuit of this, we use the uniform stability framework (Bousquet & Elisseeff, 2002; Hardt et al., 2016; Lei, 2023) of the randomized optimization algorithm.
@@ 3.4 Low-Dimensional Constraints Also Matter for Generalization
Owing to previous studies (Bousquet & Elisseeff, 2002; Hardt et al., 2016) on the generalization, $\epsilon_{\textnormal{gen}}$ is closely related to the uniform stability of the optimization algorithm.
Definition 3.12 (Uniform Stability (Bousquet & Elisseeff, 2002; Hardt et al., 2016)).
@@
Lemma 3.13 (Theorem 2.2 in Hardt et al. (2016)).
@@
Thanks to Lemma 3.13, all we have to show is that ZO-SGD under suitable assumptions is indeed uniformly stable.
Theorem 3.14 (Generalization of ZO-SGD).
@@
Suppose that (C-$1$), (C-$2$), and Assumption 3.4 hold. Further, suppose that Assumption 3.8 holds for every $\mathcal{L}_{i}$ and $\sup_{\bm{\theta},\mathbf{z}}|\mathcal{L}(\bm{\theta};\mathbf{z})|\leq 1$. Under the following parameter settings
$\alpha_{t}\leq\frac{C}{t},\quad\mu\leq\frac{1}{nL(\rho_{\textnormal{max}}+s\exp(-s/2))},$
for some fixed constant $C>0$, the generalization error bound of ZO-SGD in (3) satisfies
$\epsilon_{\textnormal{gen}}$ $\leq\mathcal{O}\Big{(}\frac{T^{1-\frac{1}{1+q}}}{n}\Big{)}$
where $q=CL\sqrt{s}$ does not rely on $d$.
@@
On the Constant $q$.  A recent study (Liu et al., 2024a) applies the uniform stability to ZO algorithms, thereby deriving generalization error bounds for various ZO gradient estimators. However, in that study, the parameter $q$ grows polynomially w.r.t. $d$, which ultimately yields a generalization error bound in the order of $\mathcal{O}(T/n)$ since $\frac{1}{1+q}\approx 0$. Moreover, the previous analysis requires $\mu=\mathcal{O}(1/nLd)$ which is tiny in extremely high-dimensional problems (ex. fine-tuning LLMs). In contrast, under our Theorem 3.14, which exploits the notion of stable rank with locally low intrinsic dimension, $q$ depends on $s$ while significantly alleviating the order of $\mu$ as $\mu=\mathcal{O}(1/nL\rho_{\text{max}})$. Consequently, we achieve much tighter bound than $\mathcal{O}(T/n)$, especially under Assumption 3.4 together with Assumption 3.8.
On Individual Assumption 3.8.  The uniform stability for generalization error analysis inherently account for single datapoint drawn at each iteration thus consider the online optimization algorithms. Therefore, the assumptions for uniform stability (ex. smoothness, Lipschitz continuity, etc) are made to hold for every per-sample loss $\mathcal{L}_{i}$ (Chen et al., 2019b; Cheng et al., 2021; Yun et al., 2022; Liu et al., 2024a). With this convention, our Assumption 3.8 is assumed for each $\mathcal{L}_{i}$. Nevertheless, we emphasize that our analysis can still yield tighter bounds independent of $d$ with only Assumption 3.4 and without relying on Assumption 3.8.
We conclude Section 3 by leaving remarks.
Remark on Section 3.  According to our theory, fast convergence requires a certain degree of overlap between the perturbation subspace and the loss landscape subspace. Hence, our result provides a necessary condition for convergence; however, the question of “how to practically choose the perturbation space” remains an open problem. Moreover, in deep learning optimization, not only is it infeasible to know the exact value of $r$, but also identifying the eigenspace of the local Hessian is even more challenging. Consequently, we aim to address the above question through an empirical study in the next section.
4 Empirical Study and Intuitions
In this section, we empirically validate the theoretical insights and explore the practical implementations of low-dimensional structures. To this end, we consider the randomized quadratic minimization problem: $\min_{\bm{\theta}\in\mathbb{R}^{d}}\frac{1}{2}\bm{\theta}^{\top}\mathbf{H}\bm{\theta}$, where $\mathbf{H}\in\mathbb{R}^{d\times d}$ is a randomly generated positive semi-definite matrix. We choose this problem because it allows direct manipulation of the Hessian structure and provides explicit access to the Hessian itself, making it straightforward to extract the principal subspace and compute the effective overlap. Moreover, this quadratic problem is widely used in prior studies (Sun & Ye, 2021; Zhang et al., 2024c) as a standard experimental framework for theoretical validation, further underscoring its suitability for our analysis.
4.1 Influence of Effective Overlap on Convergence
@@
Problem Setup. To empirically confirm that the strong impact of the effective overlap $\rho$ (Definition 3.9) on convergence, we design the following experiment. We fix the problem dimension to $d=256$ and construct a rank-$64$ Hessian whose eigenvalues decay linearly from $10$ to $1$. We then introduce a parameter $\gamma\in[0,1]$ to manually adjust $\rho$ by selecting how many of Hessian eigenvectors (specifically, the random-$\lceil s\gamma\rceil$ eigenvectors, where $s=\mathrm{srank}(\mathbf{M})$) are used. The remaining eigenvectors are randomly generated and orthogonalized against those selected, thereby controlling the effective overlap.

### v2

@@ 3.2 Convergence Analysis of Subspace ZO-SGD
    $\mathrm{intdim}\big{(}\mathbf{H}(\bm{\theta}_{t})\big{)}\leq r$ for some constant $r\leq d$.
Remark on Assumption 3.1.  This assumption extends the one originally introduced in MeZO (Malladi et al., 2023), which assumes a global $\mathcal{O}(d)$-neighborhood around $\bm{\theta}_{t}$ and takes the limit $\mu\to 0$. In contrast, our formulation allows for a strictly positive $\mu$ and only requires a local $\mathcal{O}(s)$-ball since $\mu\sqrt{d}$ is negligible under $\mu=\mathcal{O}(1/\sqrt{dT})$ (refer to Theorem 3.3). This makes our assumption significantly more practical for large-scale models where $d\gg s$, and substantially relaxes the regularity conditions on the loss landscape. When $\mathbf{H}(\bm{\theta}_{t})=\mathbf{I}_{d}$ (i.e., $r=d$), our assumption reduces to the standard $L$-smoothness condition (C-$1$). Appendix B.1 provides further discussion.
By introducing Assumption 3.1, the subspace perturbation matrix $\mathbf{M}_{t}$ and the local Hessian upper bound inevitably interact with each other. To quantify this interaction, we introduce the following key quantity for our analysis.
Definition 3.2 (Subspace Alignment).
For the matrix $\mathbf{M}_{t}$ in (2) and $\mathbf{H}(\bm{\theta}_{t})$ in Assumption 3.1, we define the subspace alignment at time $t$ as
$\rho_{t}=\frac{\mathrm{Tr}(\mathbf{M}_{t}^{\top}\mathbf{H}(\bm{\theta}_{t})\mathbf{M}_{t})}{\lambda_{\text{max}}(\mathbf{H}(\bm{\theta}_{t}))}$ (3)
@@
Remark on Subspace Alignment.  Intuitively, $\mathrm{Tr}(\mathbf{M}_{t}^{\top}\mathbf{H}(\bm{\theta}_{t})\mathbf{M}_{t})$ captures “how much the primary directions” of $\mathbf{M}_{t}$ and $\mathbf{H}(\bm{\theta}_{t})$ align each other, since both matrices are assumed to be PSD and thus act as “weights” on overlapping subspaces. Note that $0\leq\rho_{t}\leq\min\{r,s\}$ always holds.
Theorem 3.3 (Dimension-Free Rate of Subspace ZO-SGD).
@@
Novelty.  Unlike previous approaches (Rando et al., 2023b; Chen et al., 2024) that require $\mathbf{M}_{t}^{\top}\mathbf{M}_{t}=\mathbf{I}_{d}$ to ensure the unbiasedness of zeroth-order gradient estimator, our Theorem 3.3 only assumes that $\mathbf{M}_{t}$ is positive semidefinite, which introduces a slight bias while still guaranteeing a valid descent direction. Hence, our analysis can encompass a much broader range of subspace perturbation matrix $\mathbf{M}_{t}$. Additionally, we are the first to highlight the importance of subspace perturbation by introducing the concept of subspace alignment for all $\mu>0$, which sets our analysis totally apart from (Malladi et al., 2023), which focuses only on $\mu\rightarrow 0$.
@@
The convergence rate in Theorem 3.3 depends critically on the subspace alignment $\rho_{t}$, which is governed by the structure of the perturbation matrix $\mathbf{M}$. This raises a central question:
“What constitutes an effective choice of $\mathbf{M}$ for improving convergence via subspace alignment?”
The following proposition provides an intriguing preliminary insight.
Proposition 3.4 (Expected Subspace Alignment $\rho$).
Let $\mathbf{H}\in\mathbb{R}^{d\times d}$ be a fixed symmetric positive semi-definite matrix with $\lambda_{\max}(\mathbf{H})>0$. Let $\mathbf{M}\in\mathbb{R}^{d\times d}$ be a random matrix such that (i) $\mathbf{M}$ is (almost surely) an orthogonal projection matrix (i.e., $\mathbf{M}^{\top}=\mathbf{M},\mathbf{M}^{2}=\mathbf{M}$ with probability $1$), (ii) $\mathbb{E}[\mathbf{M}]=\frac{s}{d}\mathbf{I}_{d}$ holds for given $s>0$. Then, the expectation of $\rho$ over distribution of $\mathbf{M}$ is:
$\mathbb{E}[\rho]=\frac{s\operatorname{Tr}(\mathbf{H})}{d\lambda_{\max}(\mathbf{H})}.$ (4)
Remark. Proposition 3.4 shows that for any random matrix $\mathbf{M}$ satisfying these broad, non-specialized conditions, $\mathbb{E}[\rho]$ remains unchanged. Such conditions accommodate a wide variety of random matrix constructions–ranging from low-rank projections and sparse perturbations to more intricate ensembles–implying that most standard perturbation schemes share the same expected $\rho$ and, thus, exhibit comparable average convergence. This naturally raises a follow-up question: are all such choices of $\mathbf{M}$ effectively interchangeable, or do notable differences arise in practice?
To advance our investigation, we first present three canonical instantiations of $\mathbf{M}$ that fulfill the criteria specified in Proposition 3.4.
Definition 3.5 (Three Instantiations of $\mathbf{M}$).
Let $s\in\mathbb{R}$ such that $0<s\leq d$. We define three instantiations of random matrix $\mathbf{M}\in\mathbb{R}^{d\times d}$ satisfying the conditions presented in Proposition 3.4:
  1. 1.
    Low-rank Projection: $\mathbf{M}_{\mathsf{LR}}=\mathbf{U}\mathbf{U}^{\top}$, where $\mathbf{U}\in\mathbb{R}^{d\times s}$ is a random matrix with $s$ orthonormal columns. $\mathbf{U}$ is assumed to span an $s$-dimensional subspace chosen uniformly at random from the Grassmannian manifold $\mathbf{Gr}(s,\mathbb{R}^{d})$.
  2. 2.
    Sparse Perturbation: $\mathbf{M}_{\mathsf{SP}}=\operatorname*{diag}(\mathbf{m})$, where $\mathbf{m}=(m_{1},\ldots,m_{d})$ and each $m_{i}\sim\mathrm{Bernoulli}(p)$ with $p=\frac{s}{d}$.
  3. 3.
    Block Sparse Perturbation: $\mathbf{M}_{\mathsf{BSP}}=\operatorname*{diag}(\mathbf{m})$, but partition the parameters into $N$ blocks $\{B_{1},\ldots,B_{N}\}$ such that $|B_{i}|=s$ for all $i=1,\ldots,N$. For a mask $\mathbf{m}=(m_{1},\ldots,m_{d})$, we set $m_{i}=1$ if $i\in B_{j}$ where $j\sim\mathcal{U}\{1,N\}$.
Although these instantiations share the same expected $\rho$, their distributional characteristics can differ significantly, as demonstrated by the following result.
Proposition 3.6 (Upper Tail Probabilities of $\rho$).
Let $\mathbf{H}\in\mathbb{R}^{d\times d}$ be as in Proposition 3.4 and $\mathrm{intdim}(\mathbf{H})\leq r$ for some fixed $r\leq d$. Let $\widehat{\rho}$ be any threshold such that $\widehat{\rho}>\mathbb{E}[\rho]=\frac{s\mathrm{Tr}(\mathbf{H})}{d\lambda_{\max}(\mathbf{H})}$ and $\Delta=\widehat{\rho}-\mathbb{E}[\rho]>0$. Then, we have the following upper tail probabilities:
  1. 1.
    Low-Rank Projection: $\mathbb{P}(\rho\geq\widehat{\rho})\leq 2\exp\big{(}-c_{1}\frac{d\Delta^{2}}{s}\big{)}$.
  2. 2.
    Sparse Perturbation: $\mathbb{P}(\rho\geq\widehat{\rho})\leq c_{2}\frac{rs}{d\Delta^{2}}$.
  3. 3.
    Block Sparse Perturbation: $\mathbb{P}(\rho\geq\widehat{\rho})=\frac{1}{N}\big{|}\big{\{}j\in[N]:\sum_{i\in B_{j}}\mathbf{H}_{ii}\geq\lambda_{\max}\widehat{\rho}\big{\}}\big{|}=\frac{k_{\#}(\widehat{\rho})s}{d}$.
where $c_{1},c_{2}>0$ are constants and $k_{\#}(\widehat{\rho})$ counts “good” (high $\rho$) blocks.
Observations.  Proposition 3.6 highlights a clear distinction: while these methods share the same expected $\rho$, the probability distribution of $\rho$ differs considerably. In the low-rank projection and sparse perturbation cases, the likelihood of $\rho$ deviating from its mean decays exponentially and quadratically with the size of the deviation (i.e., denoted by $\Delta$), indicating a tight concentration around its expectation. In contrast, for block-sparse perturbation, although the mean remains unchanged, the probability becomes at least $1/N$ if there is any such good block.
These observations carry substantial implications. When the subspace defined by $\mathbf{M}$ is chosen purely at random, all three methods exhibit similar convergence in average. However, their distinct $\rho$ distributions become relevant once even modest adaptivity is introduced. In block sparse perturbation, for example, selectively targeting more favorable blocks (or generally identifying subspaces that yield higher $\rho$) may produce performance gains not evident from an expectation-based perspective alone.
3.3 Generalization Error Bound of Subspace ZO-SGD
All the aforementioned analyses are concerned with the convergence of the zeroth-order SGD. In this section, we aim to provide the generalization error bound of zeroth-order SGD. In pursuit of this, we use the uniform stability framework (Bousquet and Elisseeff, 2002; Hardt et al., 2016; Lei, 2023) of the randomized optimization algorithm.
@@ 3.3 Generalization Error Bound of Subspace ZO-SGD
Owing to previous studies (Bousquet and Elisseeff, 2002; Hardt et al., 2016) on the generalization, $\epsilon_{\textnormal{gen}}$ is closely related to the uniform stability of the optimization algorithm.
Definition 3.8 (Uniform Stability (Bousquet and Elisseeff, 2002; Hardt et al., 2016)).
@@
Lemma 3.9 (Theorem 2.2 in (Hardt et al., 2016)).
@@
Hence, bounding the generalization error reduces to proving the uniform stability of subspace ZO-SGD under suitable assumptions.
Theorem 3.10 (Generalization of Subspace ZO-SGD).
@@
Suppose that (C-$1$), (C-$2$), and $\mathrm{srank}(\mathbf{M}_{t})\leq s$ hold. Further, suppose that Assumption 3.1 holds for every $\mathcal{L}_{i}$ and $\sup_{\bm{\theta},\mathbf{z}}|\mathcal{L}(\bm{\theta};\mathbf{z})|\leq 1$. Under the following parameter settings $\alpha_{t}\leq\frac{C}{t}$ and $\mu\lesssim\frac{1}{nL(\rho_{\text{max}}+s\exp(-s/2))}$ for some fixed constant $C>0$, the generalization error bound of subspace ZO-SGD satisfies $\epsilon_{\textnormal{gen}}\leq\mathcal{O}\Big{(}\frac{T^{1-\frac{1}{1+q}}}{n}\Big{)}$ where $q=CL\sqrt{s}$ does not rely on $d$.
@@
On the Constant $q$.  A recent study (Liu et al., 2024a) applies the uniform stability to ZO algorithms, thereby deriving generalization error bounds for various ZO gradient estimators. However, in that study, the parameter $q=\mathcal{O}(\sqrt{d})$ grows polynomially w.r.t. $d$, which ultimately yields a generalization error bound in the order of $\mathcal{O}(T/n)$ since $\frac{1}{1+q}\approx 0$. Moreover, the previous analysis requires $\mu=\mathcal{O}(1/nLd)$ which is tiny in extremely high-dimensional spaces (ex. fine-tuning LLMs). In contrast, under Theorem 3.10, which exploits the notion of stable rank with local intrinsic dimension, $q$ depends on $s$ while significantly alleviating the order of $\mu$ as $\mu=\mathcal{O}(1/nL\rho_{\text{max}})$. Consequently, we achieve much tighter bound than $\mathcal{O}(T/n)$, especially under $s\ll d$ and Assumption 3.1.
3.4 Empirical Validation
In this section, we empirically substantiate the key theoretical insights from preceding sections using the randomized quadratic minimization problem, $\min_{\mathbf{\theta}\in\mathbb{R}^{d}}\frac{1}{2}\mathbf{\theta}^{\top}\mathbf{H}\mathbf{\theta}$ with a synthetically generated positive semi-definite Hessian $\mathbf{H}$. This tractable testbed, commonly used in prior theoretical validation studies (Sun and Ye, 2021; Zhang et al., 2024c), allows for direct verification of our claims. Details are provided in Appendix E.1.
Figure 1: Empirical validation of theoretical insights on randomized quadratic minimization. Left: Higher $\bar{\rho}$ values yield faster convergence, as predicted ($s$ is fixed). Middle: Iterations required to reach a fixed loss target decrease proportionally with $1/\bar{\rho}$. Right: Distribution of $\rho$ under three $\mathbf{M}$ types shows matching means but different concentration patterns. Horizontal blue dashed lines indicate the $\bar{\rho}$ for each method. Supplementary results are provided in Appendix F.1.
Effect of mean subspace alignment $\mathbf{\bar{\rho}}$.  Our theoretical framework (Theorem 3.3) posits a critical role for mean subspace alignment $\bar{\rho}$ in determining convergence speed. We first validate this by modulating $\bar{\rho}$ in a controlled experiment using a rank-$64$ dense Hessian with $d=256$ and a low-rank projection with $s=64$ as $\mathbf{M}$. Figure 1 (Left) confirms that higher $\bar{\rho}$ values indeed lead to faster convergence. Consistently, (Middle) demonstrates that the number of iterations required to reach a target loss (set to be the slowest one’s final loss) is inversely proportional to $\bar{\rho}$, directly corroborating our theory that the convergence rate scales with $1/\bar{\rho}$.
Distribution of subspace alignment $\mathbf{\rho}$.  Next, we investigate the practical demonstrations of Propositions 3.4 and 3.6, which concern the expected value and distributional properties of $\rho$ for different instantiations of $\mathbf{M}$. We consider three canonical instantiations provided in Definition 3.5. For this, we use a block-diagonal Hessian ($d=1024$, $16$ blocks, each rank $64$) with heterogeneous eigenspectra across blocks, a structure often used to approximate neural network Hessians (Collobert, 2004; Zhang et al., 2024b, c).
@@ 3.4 Empirical Validation
4 Revisiting Block Coordinate Descent for Zeroth-order Optimization
Our theoretical analysis showed that many subspace perturbation schemes, despite structural differences, achieve similar convergence in expectation. This intriguing finding prompts a pivotal question regarding method selection: If theory offers no clear winner, what factors matter most in practice, particularly for large-scale optimization? In such scenarios, pragmatic considerations (including memory usage and execution time) emerge as the decisive differentiators. Guided by this imperative for practical efficacy, we revisit the Block Coordinate Descent (BCD) paradigm in this section, an approach rooted in the block sparse perturbation.
4.1 Why Block Coordinate Descent?
Memory transfer is the key bottleneck in large-scale ZO.  Zeroth-order optimization, by virtue of its forward-only nature, fundamentally shares the memory-bound bottleneck commonly observed in LLM inference (Kwon et al., 2023; Kim et al., 2023; Gholami et al., 2024). For large models, the time spent loading parameters to memory can far exceed actual compute time, making memory bandwidth the dominant factor in overall latency. This challenge is even exacerbated in zeroth-order methods: a single iteration requires two forward evaluations, two parameter perturbations, and one parameter update, each involving the entire parameter. As a result, the total parameter loading per step is effectively equivalent to 5×5$\times$5 × the inference, a substantial overhead, especially for large-scale models.
BCD can reduce the memory transfer.  BCD offers a practical solution to this issue by restricting parameter perturbations and updates to a single block at each iteration. While the forward passes still involve the full model, only the selected block is loaded and modified during the perturbation and update phases. If we partition the parameters into $N$ equally sized blocks, the total parameter load per iteration is reduced to $2d+3d/N$ (two full-model forwards plus block-local perturb and update), compared to the $5d$ required by other ZO methods. This translates into a tangible reduction in memory transfer, leading to faster wall-clock training and thus improved scalability.
Importantly, BCD retains the same convergence guarantees as other qualifying subspace methods (Liu et al., 2024b; Chen et al., 2024). With its added practical efficiency, it stands out as a particularly attractive choice when theoretical distinctions are absent.
4.2 MeZO-BCD: Translating Block Sparse Perturbation into a Scalable ZO Method
Building on the practical advantages of block-wise perturbation discussed above, we now present MeZO-BCD, a concrete implementation designed to optimize large-scale LLMs under the zeroth-order paradigm. While the theoretical formulation in Section 3 abstracts perturbation as a projection onto a random subspace, MeZO-BCD operationalizes this by applying perturbation and updates to only one block per iteration.
Overall Procedure.  Let $\bm{\theta}_{t}=[\bm{\theta}_{t}^{(1)},\dots,\bm{\theta}_{t}^{(N)}]\in\mathbb{R}^{d}$ denote model parameters partitioned into $N$ disjoint blocks. Each block typically corresponds to a functional unit of LLMs, such as a transformer decoder layer (comprising attention, FFN, normalization), embedding layer, or output head (Zhang et al., 2022a; Touvron et al., 2023).
At each iteration $t$, a block index $j_{t}$ is chosen, and a local perturbation $\mathbf{u}_{t}\sim\mathcal{N}(\bm{0},\mathbf{I}_{|\bm{\theta}_{t}^{(j_{t})}|})$ is sampled using a fixed seed $s_{t}$. The block-local SPSA estimator is:
$\widehat{\mathbf{g}}_{t}=\frac{\mathcal{L}(\bm{\theta}_{t}+\mu\cdot\mathbf{e}_{B_{j_{t}}}(\mathbf{u}_{t});\mathcal{B}_{t})-\mathcal{L}(\bm{\theta}_{t}-\mu\cdot\mathbf{e}_{B_{j_{t}}}(\mathbf{u}_{t});\mathcal{B}_{t})}{2\mu}\mathbf{u}_{t},$
where $\mathbf{e}_{B_{j_{t}}}(\mathbf{u}_{t})$ embeds the block-local perturbation into the full parameter space. In practice, no actual zero-padded vector is instantiated as only the relevant block is accessed and modified.
Using the same seed $s_{t}$, we regenerate $\mathbf{u}_{t}$ for consistency during the update:
$\bm{\theta}_{t+1}^{(j_{t})}=\bm{\theta}_{t}^{(j_{t})}-\eta_{t}\cdot\widehat{\mathbf{g}}_{t},\quad\bm{\theta}_{t+1}^{(k)}=\bm{\theta}_{t}^{(k)}\text{ for }k\neq j_{t}.$
This localized update significantly reduces memory transfer overhead, as discussed above, while preserving theoretical convergence behavior.
Block selection strategy.  While the theory assumes i.i.d. uniform sampling of blocks, such sampling can lead to imbalanced short-term coverage in practice. To address this while preserving the marginal distribution $\mathbb{P}(j_{t}=j)=1/N$, we adopt a cyclic random permutation strategy, where all blocks are visited exactly once per cycle in a randomly shuffled order. This avoids starvation and maintains the theoretical convergence guarantees. Further details on MeZO-BCD are discussed in Appendix D.
4.3 Experiments on LLM fine-tuning
Table 1: Comparison of peak memory usage and iteration time across methods. Peak memory (GB) is reported on different datasets using OPT-13B, and iteration time (seconds) is measured on SST-2 across various model sizes, with speedup relative to MeZO shown in parentheses. All results are averaged over 100 training steps with a batch size of 16, using a single A100 80GB PCIe.
Method Iteration Time (s) Peak Memory (GB)
OPT-125M OPT-1.3B OPT-13B SST-2 BoolQ MultiRC